OpenStudy (anonymous):

Solve for x. Round to the nearest ten-thousandth. 36^x/2=6

jimthompson5910 (jim_thompson5910):

36^x/2=6 (6^2)^x/2=6 6^(2*(x/2)) = 6 6^x = 6 6^x = 6^1 Since the bases are equal, the exponents are equal. So x = 1

OpenStudy (anonymous):

http://www.wolframalpha.com/